# LoVR: A Benchmark for Long Video Retrieval in Multimodal Contexts

### v1

@@ 3.2.3 Full Video Captioning
$\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$ (1)
@@
3.2.4 Final Human Quality Assessment
Table 3: Baseline performance of text-to-video and text-to-clip retrieval on the LoVR benchmark.
@@ 3.2.4 Final Human Quality Assessment
CLIP 69.0 83.0 89.1 21,877,283 3,709,556 22.4 40.6 51.0 22,494,235 3,814,168
CLIP(avg. feat.) 46.0 79.0 89.0 21,877,283 3,709,556 25.7 51.6 61.4 22,494,235 3,814,168
BLIP 39.2 59.0 75.0 174,762 41,746 24.7 47.1 55.7 2,595,565 607,738
ResNet 36.3 58.0 63.0 21,096,459 3,016,545 15.7 27.4 36.0 21,691,391 3,101,613
ResNet(avg. feat.) 31.0 72.0 82.0 21,096,459 3,016,545 15.1 33.9 44.3 21,691,391 3,101,613
@@
GME-Qwen2-VL-7B 61.2 73.0 79.0 105,823,692 91,234,289 51.2 68.7 78.4 104,221,391 89,229,158
MM-Embed 58.1 71.2 78.0 103,957,148 86,354,225 47.6 67.0 75.8 106,888,800 88,789,465
MM-Embed(avg. feat.) 52.0 79.0 89.0 103,957,148 86,354,225 43.8 70.1 80.1 106,888,800 88,789,465
jina-clip-v2 57.0 69.3 76.0 92,332,391 83,296,271 45.7 65.1 73.0 94,329,192 86,815,279
3.3 Dataset Statistics
As shown in Table 1, the final dataset of LoVR comprises a total of 467 videos and 40,804 fine-grained clips. Compared to previous benchmarks such as MSR-VTT xu2016msr , MSVD chen2011collecting , and VDC chai2024auroracap , our benchmark offers a significantly larger dataset, longer video durations, and more extensive caption lengths, considering both clip-level and video-level annotations. Annotating with VLM+Human achieves higher scalability than human annotation and better quality than pure VLM-based methods. Furthermore, our benchmark highlights the significance of thematic information, which is neglected by many existing benchmarks. The dataset is divided into a training set and a test set, each containing both long video and clip data. Specifically, to ensure accurate correspondence between clips and captions, we construct the test set by randomly selecting 100 videos to serve as the video portion of the test set. All the clips of these videos form the clip portion of the test set. The remaining 367 videos and all of their associated clips constitute the training set.
4 Experiments
4.1 Experimental Settings
4.1.1 Evaluation Metrics.
@@ 4.1 Experimental Settings
Retrieval Accuracy.
This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results. This metric reflects the recall capability of the retrieval model.
Retrieval Time.
@@
4.1.2 Baselines
For long video-text retrieval, the encoder plays a central role in modeling visual and text content. Therefore, we select three categories of widely used encoders for evaluating our benchmark. The following are the baselines used:
Image Encoders.
BLIP li2022blip uses an encoder-decoder architecture for vision-language understanding. ResNet50 (CLIP) he2016deep ; radford2021learning is a CNN-based image encoder with skip connections, pre-trained within the CLIP framework. ViT-L/14 (CLIP) dosovitskiy2020image ; radford2021learning is a Vision Transformer trained using the CLIP objective for global image representation.
Video Encoders.
VideoCLIP xu2021videoclip extends CLIP to video-text contrastive learning using temporal aggregation. VideoCLIP-XL is a long-video understanding model designed with high-resolution input to enhance comprehension. VideoCLIP-XL-v2 is an enhanced version of VideoCLIP-XL with better modeling capabilities. wang2024videoclip
Multimodal Embedding Models.
GME-Qwen2-VL-7B-Instruct zhang2024gme is a multilingual multimodal embedding model based on Qwen2-VL. MM-Embedlin2024mm is a universal multimodal retrieval model that integrates multimodal large language capabilities with a modality-aware negative mining strategy. jina-clip-v2 koukounas2024jina extends the CLIP architecture to support multiple languages, high-resolution image inputs, and compact embedding representations using multi-resolution learning.
Table 4: Baseline performance of video-to-text and clip-to-text retrieval on the LoVR benchmark.
@@ 4.1.2 Baselines
CLIP(avg. feat.) 19.0 55.0 71.0 21,877,283 3,709,556 27.6 61.7 78.0 22,494,235 3,814,168
BLIP 37.0 60.0 67.0 167,839 18,876 49.4 76.8 85.2 482,842 244,694
ResNet(avg. feat.) 12.0 35.0 60.0 21,096,459 3,016,545 28.9 54.0 66.8 21,691,391 3,101,613
@@
GME-Qwen2-VL-7B 57.2 83.8 90.2 105,823,692 91,234,289 75.2 93.9 95.1 104,221,391 89,229,158
MM-Embed(avg. feat.) 56.0 84.0 91.0 103,957,148 86,354,225 77.7 94.9 97.6 106,888,800 88,789,465
jina-clip-v2 56.3 83.2 89.4 92,332,391 83,296,271 75.3 92.6 93.4 94,329,192 86,815,279
4.2 Baseline Performance on Video Retrieval and Clip Retrieval
We evaluate a range of baseline methods on the LoVR benchmark across two core tasks: video retrieval and clip retrieval, in both text-to-video/clip and video/clip-to-text directions (see Tables3, and 4). It can be seen that even the top-performing method, MM-Embed (avg. feat.), achieves only 56.0% Recall@1 in video-to-text and 77.7% Recall@1 in clip-to-text retrieval, indicating the difficulty of our benchmark. While models like VideoClip-XL-v2 perform reasonably well on short clips (e.g., 76.0% Recall@1 in clip-to-text), their effectiveness drops notably on full-length videos (59.0% Recall@1). Similarly, image-based encoders such as CLIP show limited generalization to long-form content, with a maximum text-to-video Recall@1 of 69.0%, which is a relatively low score given the complexity of LoVR’s real-world scenes. These findings underscore the unique challenges of the LoVR benchmark, including long temporal dependencies, semantic diversity, and high alignment requirements between modalities, thus providing a meaningful and realistic testbed for advancing robust, scalable long video-text retrieval systems.
4.3 Baseline Performance on Theme Retrieval
As shown in Table 5, we evaluate the performance of baseline models on theme-to-clip and clip-to-theme retrieval tasks. Here, ‘theme’ refers to the thematic information of each clip, focusing on high-level semantic aspects such as narrative themes, emotional tones, and contextual settings.
Considering the ambiguity in thematic information, we expand the evaluation to include Recall@25 and Recall@50. Despite this expansion, the overall performance remains suboptimal, with the highest theme-to-clip Recall@50 score reaching only 77.5%. This suggests that current models still struggle to effectively capture and represent thematic characteristics. Notably, clip-to-theme retrieval performs slightly better than the reverse task. One reason is that the number of distinct themes per clip is relatively limited, which inherently reduces the complexity of the retrieval task.
Table 5: Baseline performance of theme-to-clip and clip-to-theme retrieval. Avg. feat. denotes the feature of the video is the average frame feature.
Model Theme-to-Clip Clip-to-Theme
R@25 R@50 R@25 R@50
CLIP(avg. feat.) 34.3 41.9 43.5 57.2
BLIP 68.1 77.5 61.1 70.5
ResNet 14.9 21.3 - -
ResNet(avg. feat.) 23.9 30.9 42.7 58.0
VideoClip 39.7 52.9 73.2 82.1
VideoClip-XL 45.3 56.1 79.8 86.4
VideoClip-XL-v2 49.3 58.2 80.0 85.6
GME-Qwen2-VL-7B 43.9 49.8 74.6 83.9
MM-Embed 38.2 45.6 - -
MM-Embed(avg. feat.) 47.4 55.4 75.3 85.5
jina-clip-v2 38.3 45.1 73.1 82.4
@@ 4.3 Baseline Performance on Theme Retrieval
Figure 5: Case Study of LoVR. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
4.4 Case Study
@@
As shown in Figure 5, two example failure cases are presented. In Query 1, due to the length of the input text, we omit the content of the middle portion. It can be observed that under long-query conditions, the model fails to process the entire context, resulting in the neglect of information at the end of the query. This leads to the omission of critical details — for instance, the key element "pan" in this case is ignored. In Query 2, the model demonstrates limitations in capturing theme and atmosphere information. Although four videos related to urban city scenes are retrieved, only the first one accurately reflects the required ‘solemn and sorrowful atmosphere.’ The remaining results fail to incorporate this essential emotional aspect, leading to incorrect retrievals. These cases highlight the importance of thematic and affective information in video understanding and retrieval tasks.
5 Conclusion
@@ Appendix A Details of LoVR Construction
A.1 Implementation Details.
VLM settings
VLMs utilized for caption generation include Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-7B-Instruct bai2025qwen2 , and Qwen2-VL-72B-Instruct wang2024qwen2 . The initial temperature is set to 0.9 during caption generation.
Experiment settings
@@
The inter-frame difference scores shown in Figure 3 are calculated by computing the differences between the average pixel values of consecutive frames. To ensure a robust evaluation, we run each quantitative experiment three times and take the average. In the baseline approach, all encoder models first vectorize the videos and the query captions into embeddings. Then, the cosine similarity between each video feature and the query caption feature is computed and used for ranking. The models can be categorized into two groups based on their architecture and processing capabilities:
  • Image encoders and multimodal models that can not process video: These models include CLIP, ResNet, and MM-Embed, which are originally designed for image-level representation and require adaptation to handle video inputs. During the experiments, we set the number of extracted frames used by these models to 10. Two strategies are commonly applied:
    1. Frame-wise maximum similarity: Frames are sampled from each video, and individual frame embeddings are extracted. The cosine similarity between each frame and the query caption is calculated, and the highest score across all frames is taken as the overall similarity between the video and the caption.
    2. Average feature pooling (‘avg. feat.’): Similar to the above, but instead of selecting the maximum similarity, the frame embeddings are averaged to form a single video-level embedding, which is then used to compute similarity with the caption. This strategy is explicitly marked with the suffix ‘avg. feat.’ in Tables 3, 4, and 5. Models without this suffix follow the first strategy.
  • Video encoders and multimodal models that can process video: These models include BLIP, VideoClip, VideoClip-XL, GME, and jina-clip, which are inherently capable of processing temporal information in videos. They directly encode the entire video into a single video-level embedding, based on which the similarity with the text query is computed, without requiring manual aggregation or sampling.
A.2 Computational Costs.
@@
The caption generation process using Qwen2.5-VL-72B-Instruct bai2025qwen2 required approximately 720 GPU hours. Evaluating the generated captions with the EVQAScore liang2024evqascore model added approximately 100 GPU hours to the total computational load. Consequently, the construction of LoVR incurred a total computational cost of roughly 820 GPU hours, based on NVIDIA H800 GPU usage.
Appendix B Extra Experiment Analysis
B.1 Duration of Clips
As shown in the left plot of Figure6, the video duration distribution in our dataset exhibits a balanced coverage of both short and medium-length videos, distinguishing it from existing benchmarks like MSVD chen2011collecting , MSR-VTT xu2016msr and TGIF tgif-cvpr2016 that primarily focus on short clips. Our dataset includes a significant number of medium-length videos (40-120 seconds) that account for approximately 15% of the total collection. This distribution schema enables comprehensive evaluation of video retrieval models, testing their capability to handle both brief visual semantics in short clips and complex temporal relationships in longer videos.
Figure 6: Left: distribution of clip durations in the LoVR dataset. Right: comparison of Recall@K performance among baseline models.
B.2 Comparative Analysis of Recall@K for Baseline Models
The Recall@K curves(right portion of Figure 6)across different models at the clip-level reveal notable performance differences, reflecting variations in model architecture, training strategies, and semantic understanding capabilities. Specifically, the VideoClip-XL series demonstrates the best overall performance, with VideoClip-XL-v2 achieving a Recall@1 of 56.5% and a Recall@20 of 91.5%, significantly outperforming other models. This indicates its stronger ability in fine-grained video understanding. In contrast, early-stage models such as CLIP and ResNet exhibit relatively low accuracy even at small k values, with slower growth trends, highlighting their limitations in capturing complex video semantics.
However, despite some models showing promising results at higher k values, most models still fall short in achieving satisfactory performance on metrics like Recall@1 or Recall@5. This suggests that there is substantial room for improvement, particularly in modeling semantic details and dynamic scene transitions. These results further confirm the high level of challenge posed by our benchmark, which effectively discriminates between model capabilities and provides a meaningful testbed for future research in video-language understanding.
B.3 Impact of Frame Count in Video Embedding on Retrieval Performance
Table 6: Baseline performance of text-to-video and video-to-text retrieval under different frame counts.
Model Frame Count Text-to-Video Retrieval Video-to-Text Retrieval
R@1 R@5 R@10 R@1 R@5 R@10
ResNet 10 32.0 72.0 81.0 12.0 35.0 59.0
20 31.0 72.0 81.0 12.0 35.0 59.0
Clip 10 46.0 78.0 89.0 19.0 55.0 71.0
20 46.0 79.0 89.0 19.0 54.0 71.0
Table 7: Baseline performance of text-to-clip and clip-to-text retrieval under different frame counts.
Model Frame Count Text-to-Clip Retrieval Clip-to-Text Retrieval
R@1 R@5 R@10 R@1 R@5 R@10
ResNet 10 15.1 34.5 44.4 28.8 54.3 66.6
20 15.2 34.9 43.6 29.0 54.0 66.1
Clip 10 25.8 51.6 61.2 27.4 61.9 77.8
20 25.8 51.1 61.4 27.5 61.2 77.7
From Tables 6, 7, and 8, we can observe that the retrieval performance varies under different frame counts, indicating that the number of video frames has an impact on the effectiveness of multimodal retrieval tasks. Although there are slight fluctuations in some metrics, the overall trend suggests that incorporating more frames helps improve the model’s ability to understand and represent video content, thereby enhancing cross-modal matching accuracy. Since videos are composed of multiple frames, a higher frame count generally means more comprehensive visual information. However, using more frames also entails greater computational costs and longer retrieval times, which requires a trade-off based on practical considerations.
At the same time, the results across all tasks reaffirm the high level of challenge posed by the LoVR benchmark dataset. Even when using well-established visual feature extractors such as ResNet, the overall retrieval performance still leaves significant room for improvement, particularly in clip-level retrieval tasks. This highlights the need for stronger cross-modal alignment capabilities in the models.
Table 8: Baseline performance of theme-to-clip and clip-to-theme retrieval under different frame counts.
Model Frame Count Theme-to-Video Retrieval Video-to-Theme Retrieval
R@25 R@50 R@25 R@50
ResNet 10 24.6 31.8 42.9 57.5
20 24.0 31.3 42.7 57.8
Clip 10 34.7 42.1 43.6 57.4
20 35.6 42.4 43.5 57.7
Appendix C Details of Human Quality Assessment

### v2

@@ 3.2.3. Full Video Captioning
(1) $\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$
@@
Table 2. Baseline performance of text-to-video and text-to-clip retrieval on the LoVR benchmark.
@@
CLIP 23.34 43.90 54.39 2544747 1431709 18.78 37.28 46.31 222270298 125052299
MetaCLIP-ViT-H-14 26.76 53.10 65.52 1165648 313637 25.26 46.10 55.09 101813266 27394590
siglip-base-patch16-224 19.06 37.90 52.25 1376746 344674 19.34 36.81 45.26 120251547 30105548
siglip2-base-patch16-224 10.06 27.41 37.04 1336867 269687 8.35 19.03 25.28 116768361 23555767
EVA02-CLIP-B-16 26.34 51.61 62.96 7491183 5156938 23.37 44.16 53.54 654315581 450431558
PE-Core-B16-224 23.34 48.60 58.24 7460584 4967311 19.95 38.42 47.30 651642897 433868639
@@
GME-Qwen2-VL 19.91 41.54 53.32 19242308 18301540 53.53 75.89 82.72 1680714676 1598543574
$M$-Embed 14.13 31.05 40.69 2030535 1168405 32.32 52.84 60.96 177356626 102054123
LanguageBind-Video 42.61 66.60 77.94 180170 180170 40.30 65.13 74.26 15742371 15742370
3.2.4. Final Human Quality Assessment
3.3. Dataset Statistics
As shown in Table 1, the final dataset of LoVR consists of 467 videos and 40,804 fine-grained clips. Compared to conventional benchmarks such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011), our benchmark provides a larger scale of data, longer video durations, and more comprehensive captions at both the clip level and video level.
The annotation process employs a "VLM + Human" strategy, which achieves greater scalability than pure human annotation and higher quality than purely VLM-based methods. Furthermore, in the prompts designed for VLM annotation, the model is explicitly instructed to identify thematic information of the video, serving as contextual reference for video content. The entire dataset is designated as the test set, and all data are used during experimental evaluation.
Figure 5 illustrates key characteristics of the dataset. The figure (b) shows that the video captions in our dataset are notably longer and contain more detailed information. The figure (c) presents the duration distribution of clip-level videos. As the fundamental units of long videos, these clips cover short to medium lengths, and their combination forms the long video structure. This contrasts with existing benchmarks like MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and TGIF (Li et al., 2016), which primarily focus on short video clips. A considerable proportion of medium-length videos (40–120 seconds) is included, accounting for approximately 15% of the total collection. This distribution allows for a comprehensive evaluation of video retrieval models, testing their ability to comprehend brief visual semantics in short clips as well as complex temporal relationships in longer videos. The figure (d) from the left displays the distribution of clip caption lengths. It can be observed that the captions are highly detailed, with the majority exceeding 1,000 tokens, offering rich descriptions of the clip content.
4. Experiments
4.1. Experimental Settings
4.1.1. Evaluation Metrics.
@@ 4.1. Experimental Settings
Retrieval Accuracy
This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results and reflects the recall capability of the retrieval model. For video retrieval with image-based models, we adopt the average feature pooling strategy, where frame-level embeddings are averaged into a single video-level representation before computing similarity with the caption.
Retrieval Time
@@
Table 3. Baseline performance of video-to-text and clip-to-text retrieval on the LoVR benchmark.
@@
CLIP 16.70 35.55 47.11 2544747 1431709 18.29 35.99 44.76 222270298 125052299
MetaCLIP-ViT-H-14 22.27 46.68 58.89 1165648 313637 28.32 51.02 60.52 101813266 27394590
siglip-base-patch16-224 15.84 36.40 45.61 1376746 344674 21.02 40.23 49.50 120251547 30105548
siglip2-base-patch16-224 7.49 23.55 31.48 1336867 269687 8.21 18.51 24.62 116768361 23555767
EVA02-CLIP-B-16 16.27 38.97 48.61 7491183 5156938 23.25 44.05 53.83 654315581 450431558
PE-Core-B16-224 18.42 35.33 43.68 7460584 4967311 18.43 36.00 44.58 651642897 433868639
@@
GME-Qwen2-VL 8.35 20.13 28.91 1680714676 1598543574 42.82 66.03 74.15 19242308 18301540
$M$-Embed 14.99 37.26 48.82 2030535 1168405 32.81 53.55 61.75 177356626 102054123
LanguageBind-Video 37.47 60.39 73.23 180170 180170 35.83 61.12 70.84 15742371 15742370
4.1.2. Baselines
In order to fully evaluate our benchmark, we select three categories of widely used retrieval models. The following are the baselines used:
Image Encoders.
CLIP (Radford et al., 2021) is a foundational vision-language model that learns a shared embedding space for images and text through contrastive pre-training. MetaCLIP-ViT-H-14 (Xu et al., 2023) is a large-scale CLIP variant trained on a vast dataset, utilizing a Vision Transformer (ViT-H/14) backbone for enhanced visual representation. SigLIP-base-patch16-224 (Zhai et al., 2023) improves upon CLIP by employing a sigmoid-based loss function that operates on individual image-text pairs. SigLIP2-base-patch16-224 (Tschannen et al., 2025) is an advanced iteration of SigLIP that integrates multiple pretraining techniques into a unified recipe. EVA02-CLIP-B-16 (Fang et al., 2024) leverages the robustly optimized EVA-02 Vision Transformer as its image encoder, offering strong performance in vision-language tasks. PE-Core-B16-224 (Bolya et al., 2025) is a competitive perception encoder trained with a robust vision-language learning objective, designed for high performance in image and video understanding tasks.
Video Encoders.
VideoCLIP-XL (Wang et al., 2024b) is specifically designed for long video understanding, leveraging a large-scale dataset of video-long description pairs to enhance its comprehension of extended temporal content. VideoCLIP-XL-v2 (Wang et al., 2024b) is an optimized successor to VideoCLIP-XL, trained with additional data and refined techniques to further boost its capabilities in long-form video-text retrieval.
Multimodal Embedding Models.
GME-Qwen2-VL (Zhang et al., 2024b) is a unified multimodal embedding model built upon the Qwen2-VL multimodal large language model, capable of generating joint embeddings for text, images, and image-text pairs. MM-Embed (Lin et al., 2024) is a universal multimodal retrieval model based on a multimodal large language model (MLLM) with a bi-encoder architecture, designed for efficient and broad-spectrum cross-modal search. LanguageBind-Video (Zhu et al., 2023c) adopts a language-centric approach to align video features with other modalities in a shared semantic space.
4.2. Baseline Performance on Video Retrieval and Clip Retrieval
We benchmark a comprehensive suite of models on the LoVR dataset across two fundamental retrieval tasks: text-to-video/clip and video/clip-to-text (see Tables 2 and 3). Overall, LanguageBind-Video emerges as the most effective model among all evaluated baselines, achieving the highest performance in both text-to-video (42.6% R@1) and video-to-text (37.5% R@1) retrieval. Its clip-level performance is also strong, reaching 40.3% R@1 in text-to-clip and 35.8% R@1 in clip-to-text retrieval. In contrast, image-based encoders such as CLIP and MetaCLIP-ViT-H-14 show limited scalability when handling long-form videos. For instance, CLIP achieves only 23.3% R@1 in text-to-video retrieval and 16.7% R@1 in video-to-text retrieval, demonstrating poor temporal modeling capacity. VideoClip-XL-v2, as a dedicated video encoder, exhibits solid performance in short-clip retrieval (55.3% R@1 in text-to-clip and 48.5% R@1 in clip-to-text), but its performance on full-length video retrieval remains moderate (29.9% R@1 in text-to-video and 25.1% R@1 in video-to-text). Among multimodal embedding models, GME-Qwen2-VL and MM-Embed demonstrate good clip-level recall (53.5% and 32.3% R@1 in text-to-clip, respectively), yet both suffer a substantial drop on long videos, indicating that multimodal embeddings trained primarily on images struggle to generalize to extended temporal contexts.
These results collectively highlight the inherent difficulty of the LoVR benchmark. Its long temporal dependencies, rich semantic variation, and stringent cross-modal alignment requirements make it a rigorous and realistic testbed for evaluating scalable long video–text retrieval systems.
Figure 5. Recall@K performance comparison across baseline models on LoVR clip-level retrieval.
Table 4. Long video retrieval performance under different frame counts across retrieval tasks.
Model Frames Text-to-Clip Text-to-Video Clip-to-Text Video-to-Text
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Clip 10 18.82 37.50 46.29 23.34 43.90 54.39 18.16 35.82 44.64 16.49 35.76 46.90
20 18.94 37.50 46.49 23.13 43.90 54.39 18.20 35.82 44.75 16.49 35.76 46.68
30 18.78 37.28 46.31 23.34 43.90 54.39 18.20 35.99 44.76 16.70 35.55 47.11
VideoCLIP-XL 10 49.06 73.69 81.66 30.62 54.60 64.45 45.36 70.96 79.41 30.84 54.39 66.81
20 48.31 73.34 81.35 30.84 55.46 67.02 44.66 70.67 78.86 29.76 54.60 66.17
30 47.33 72.67 80.81 30.84 54.82 65.74 44.18 69.95 78.64 29.12 54.39 66.38
$M$-Embed 10 32.64 53.21 61.35 14.13 31.69 40.90 32.95 53.69 61.80 14.99 37.47 49.25
20 32.36 53.04 61.20 14.13 31.69 40.69 32.81 53.71 61.76 14.99 37.47 49.25
30 32.32 52.84 60.96 14.13 31.05 40.69 32.81 53.55 61.75 14.99 37.36 48.82
4.3. Comparative Analysis of Recall@K for Baseline Models
Figure 5 presents the Recall@K performance of representative baselines on the clip-level retrieval task. VideoClip-XL achieves the strongest overall performance, with Recall@1 of 47.3% and Recall@20 of 87.5%, demonstrating superior temporal modeling and cross-modal alignment. LanguageBind follows closely, reaching 40.3% at Recall@1 and 81.9% at Recall@20, while MM-Embed remains moderate, with Recall@1 of 32.3% and Recall@20 below 69.0%. In contrast, CLIP performs weakest, starting at only 18.8% and reaching just 55.9% even when $K=20$.
Notably, despite steady improvement across all models as $K$ increases, even the best-performing model fails to achieve perfect recall at large $K$. This plateau effect underscores the intrinsic difficulty of the LoVR benchmark: it requires precise long-range temporal reasoning and robust semantic grounding to retrieve relevant content among diverse, real-world video scenes. Consequently, LoVR serves as a rigorous and discriminative benchmark for advancing the next generation of video–language retrieval systems.
@@ 4.3. Comparative Analysis of Recall@K for Baseline Models
Figure 6. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
4.4. Impact of Frame Count in Video Embedding on Retrieval Performance
As shown in Table 4, varying the number of sampled frames leads to only marginal differences in retrieval accuracy. While using more frames provides additional visual information, the improvements are inconsistent and often negligible across tasks. For example, CLIP shows virtually no improvement when increasing frames from 10 to 30 across all tasks, with performance remaining nearly flat. MM-Embed exhibits similar behavior with minimal variations. Even VideoCLIP-XL—though achieving the best overall performance—demonstrates mixed results: while Text-to-Video tasks show slight improvements with more frames, Text-to-Clip and Video-to-Text tasks actually perform best with 10 frames, suggesting that simply adding more frames may introduce noise rather than useful information.These results underscore the inherent difficulty of the LoVR benchmark. Despite employing advanced visual feature extractors, the retrieval performance remains far from satisfactory, particularly evident in the large performance gaps between different model architectures. The benchmark therefore presents a demanding setting that cannot be solved by naive frame sampling strategies alone, but instead calls for more sophisticated algorithms and models capable of robust long-video understanding and cross-modal alignment.
4.5. Case Study
@@
As shown in Figure 6, two example failure cases are presented. In Query 1, due to the length of the input text, we omit the content of the middle portion. It can be observed that under long-query conditions, the model fails to process the entire context, resulting in the neglect of information at the end of the query. This leads to the omission of critical details — for instance, the key element "pan" in this case is ignored. In Query 2, the model demonstrates limitations in capturing theme and atmosphere information. Although four videos related to urban city scenes are retrieved, only the first one accurately reflects the required ‘solemn and sorrowful atmosphere.’ The remaining results fail to incorporate this essential emotional aspect, leading to incorrect retrievals. These cases highlight the importance of accurate long context understanding in long video retrieval tasks.
5. Conclusion
@@
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
@@ Appendix A Details of LoVR Construction
A.1. Implementation Details.
@@
The inter-frame difference scores shown in Figure 3 are calculated by computing the differences between the average pixel values of consecutive frames. To ensure a robust evaluation, we run each quantitative experiment three times and take the average. In the baseline approach, all encoder models first vectorize the videos and the query captions into embeddings. Then, the cosine similarity between each video feature and the query caption feature is computed and used for ranking.
A.2. Computational Costs.
@@
The caption generation process using Qwen2.5-VL-72B-Instruct (Bai et al., 2025) requires approximately 720 GPU hours. Evaluation of the generated captions with the EVQAScore (Liang et al., 2024) model adds approximately 100 GPU hours to the total computational load. Consequently, constructing LoVR incurs a total computational cost of roughly 820 GPU hours on NVIDIA H800 GPUs. All experiments are conducted on NVIDIA H800 GPUs.
Appendix B Details of Human Quality Assessment